1/1

Heinrich Heine,  Heinrich Heine,
Lurleia  The Lorelei
tradukita de N. N. 01  tradukita de Mark Twain

An ancient legend of the Rhine

Ignoro, quid id sibi velit, Tristissimus cur sim, Antiqui aevi fabellam Cur saepe volverim.

I cannot divine what it meaneth, This haunting nameless pain: A tale of the bygone ages Keeps brooding through my brain:

Vesperascit et frigescit, Et Rhenus leniter it, Cacumen montis lucescit, Dum Phoebus occidit.

The faint air cools in the gloaming, And peaceful flows the Rhine, The thirsty summits are drinking The sunset’s flooding wine;

Sedet in summo montis Virgo pulcherrima, Auro nitet gemma frontis, Se pectit auricoma.

The loveliest maiden is sitting High-throned in yon blue air, Her golden jewels are shining, She combs her golden hair;

Aureolo pectine pectit, Carmen canens procul, Mirandum id habet modum Nec non virilem simul.

She combs with comb that is golden, And sings a weird refrain That steeps in a deadly enchantment The listener’s ravished brain:

In cymba navitam mille Angores feri tenent, Non videt scopulos ille, Ocli non si sursum vident.

The doomed in his drifting shallop, Is tranced with the sad sweet tone, He sees not the yawing breakers, He sees but the maid alone:

Opinor undas devorare Nautam cum navicula, Effecit solo canendo Lurleia id dea.

The pitiless billwos engulf him!- So perish sailor and bark; And this, with her baleful singing, Is the Lorelei’s gruesome work.

Traduko de la Germana poemo “Die Lorelei” de Heinrich Heine (∗1797-12-13 – †1856-02-17) en la Latinan de N. N. 01.

Arg-2-179 (2010-02-11 14:00:52)

Traduko de la Germana poemo “Die Lorelei” de Heinrich Heine (∗1797-12-13 – †1856-02-17) en la Anglan de Mark Twain.

Arg-2-11 (2003-10-13 04:42:59)

A Tramp Abroad. Vol 1-2. Leibzig: Tauchnitz, 1880 Band I, Mark Twain 1880 (rf. http: // www. loreley. com/ loreley/ marctwai. htm )

http://www.poezio.net/version?poem-id=2&version-id=179,11